# Diagnosing and Mitigating Semantic Inconsistencies in Wikidata’s Classification Hierarchy

### v1

@@ Abstract
1 Introduction
Established in 2012, Wikidata has evolved into one of the internet’s most prominent sources of structured information, encompassing over 120 million entities [8]. Its success stems from a unique model that combines community contributions with data integration from diverse external sources like Gene Ontology and Wikipedia [7]. This openness has positioned Wikidata as a central resource for knowledge graph research, yet its relatively loose editorial policy introduces significant challenges to its taxonomic integrity and data quality.
While numerous quality issues exist, such as property constraint violations and entity duplication [7], a particularly fundamental problem lies within its classification hierarchy. The predicates P31 (instance of) and P279 (subclass of) form the taxonomic backbone of the knowledge graph, connecting over 99% of its entities. However, the misuse and conflation of these two critical relationships have led to what Dadalto et al. (2022) described as a ”large-scale conceptual disarray” [2]. Their work revealed frequent anti-patterns where entities are simultaneously treated as both instances and classes, undermining the logical consistency of the classification system.
The consequences of such taxonomic inconsistencies are far-reaching. A clear distinction between instances and classes is crucial for the performance of downstream applications, including knowledge graph embedding, where it enables more precise modeling of relational patterns [6]. Furthermore, structural issues like redundant connections and classification errors can introduce noise that degrades the efficiency and accuracy of reasoning tasks [9]. For a large, multi-domain knowledge graph like Wikidata, maintaining internal structural coherence is therefore essential for its reliability and utility.
@@ 1 Introduction
2 Related Work
2.1 Wikidata Quality
@@
Wikidata’s status as a premier large-scale open knowledge graph is built upon its principles of community control, plurality, and reliance on secondary data[8], which have fueled its expansion to over 120 million entities by 2025. However, this scale and openness also introduce significant data quality challenges. A comprehensive study by Shenoy et al. (2021)[7] identified a spectrum of quality issues, including property constraint violations, entity duplication, and inconsistent naming conventions. Among these, the conceptual confusion between classes (types) and instances (individuals) stands out as a fundamental and pervasive problem.
Dadalto et al. (2022) [2]provide compelling evidence of this ”large-scale conceptual disarray” within Wikidata’s multi-level taxonomies. Through an analysis of anti-patterns involving the
@@
P31 (instance of) and P279 (subclass of) relations, their work reveals that a vast number of entities are simultaneously treated as both types and instances. This issue is particularly acute in domains such as biology and biochemistry, where concepts like gene and protein are frequently misrepresented. The authors attribute this problem to several root causes: the inherent ambiguity of natural language constructs like ”is-a” (Brachman, 1983[1]), the multi-faceted nature of many real-world concepts, and practical issues such as large-scale automated data imports and the limited adoption of formal classification schemes. These findings underscore the critical need for robust methods to detect and resolve such taxonomic inconsistencies, which is a primary motivation for our current research.”
2.2 Redundant Connections and KG-based tasks
@@
In the context of large-scale knowledge graphs such as Wikidata, the issue of edge redundancy presents a nuanced challenge with both detrimental and beneficial aspects. From a data quality and maintenance perspective, redundancy is traditionally viewed as a negative attribute. Foundational database theories, as discussed by Ferrarotti et al. (2009)[4], highlight that redundant relations can lead to update anomalies and data inconsistencies. This perspective is echoed in knowledge graph construction, where methodologies often aim to build low-redundant and high-accuracy graphs, treating redundancy as noise that compromises data integrity (Li et al., 2023[5]). Furthermore, formal validation approaches, such as formalizing Wikidata’s property constraints using SHACL and SPARQL, implicitly target redundancy, as redundant edges that are not properly maintained can lead to constraint violations and complicate data validation efforts (Ferranti et al., 2024[3]).
@@
The negative impact of redundancy is particularly pronounced in downstream knowledge graph reasoning tasks. Research by Wang et al. (2023)[9] demonstrates that path redundancy in Graph Neural Network (GNN)-based models can increase transformation error entropy, thereby degrading both the efficiency and accuracy of the reasoning process. Their proposed solution involves a mechanism to prune redundant paths during reasoning, underscoring the performance cost associated with such redundancy.
@@
However, a complete removal of all logically redundant edges may not be the optimal strategy. The utility of a knowledge edge is often context-dependent. Yan et al. (2020)[10] propose a model that learns to generate and filter contextualized knowledge structures for specific commonsense reasoning tasks. Their work suggests that an explicitly stated edge, even if logically inferable and thus ”redundant”, can serve as a powerful and direct signal or a ”shortcut” for reasoning. Eliminating such edges globally could remove valuable heuristics and harm the performance on certain tasks. This indicates that a more sophisticated approach than outright deletion is required. The existing literature thus points towards the need for a filtering mechanism, rather than complete removal, to manage redundancy—a perspective that motivates our present work.
3 Methodology
@@
In our ongoing research, we conducted a three-stage analysis of the overall data structure of Wikidata to identify and mitigate semantic ambiguities across its classification system.
@@
To support this, we constructed a graph-structured representation of Wikidata using the complete set of P31 and P279 triples. Based on the connectivity patterns among parent classes, intra-graph distances, and hierarchy topology, we defined four dimensions of risk to assess the consistency of each entity’s classification. The resulting scores were integrated into a web-based user interface that enables real-time inspection of the semantic risk associated with any given entity.
@@
3.1 CME Detection Pipeline
In Wikidata, the mixed use of P31 and P279 can lead to confusion in classification hierarchy and meaning. The Figure1 shows some typical errors, as well as the link forms we consider correct in this study. An entity that is linked to a higher level through the P279 relationship should be considered a ”class” entity. As a class entity, it can be an instance of the ”class” (abstract definition) in Wikidata, but it should not be an instance of other entities with actual meaning. Conversely, if an entity is already linked to a meaningful class entity through P31, it should itself be an instance rather than a subclass, and it should not have P279 relationships from other entities. The ’complete and correct’ patterns depicted in Figure 1 are derived from foundational principles of multi-level modeling theory[2], which enforce a strict stratification between instances (0-order entities) and classes (first-order, second-order, etc.). Any deviation from these patterns, such as an entity being both an instance and a subclass, violates these principles.
Fig. 1: Error Types in P31/P279 Relations
@@ 3.1 CME Detection Pipeline
We tested this method across six selected knowledge domains (Organism, Music, Human, Gene, Geography, Color). Each domain was sampled five times with 10,000 entities per sample. For each, we randomly selected 100 entities for manual validation.The result will be showed in the Results part.
3.2 Entity-Level Semantic Risk Evaluation Metrics
@@
To assess the degree of semantic ambiguity or multi-meaning risk associated with individual Wikidata entities, we designed a scoring system based on the following four dimensions:
  1. 1.
    Number of P31 and P279 links: Prior statistical studies show that the average number of P31 relations per entity is approximately 1.3, while the average number of P279 relations is around 1.2. An unusually high count of either predicate may indicate that the entity plays multiple roles or is mapped to overlapping semantic categories.
  2. 2.
    Distance among parent classes: We compute the pairwise distances between all direct parent classes of an entity within the P31–P279 subgraph. If these distances are large, it implies that the associated parent classes originate from semantically distant regions of the knowledge graph, thus increasing the likelihood of conceptual inconsistency.
  3. 3.
    Variance of depth to root: For each parent class, we calculate its shortest path to the root node (typically Q35120). The variance of these distances reflects the structural stability of the entity’s classification. A high variance indicates uncertainty or inconsistency in taxonomic hierarchy—i.e., whether the entity’s parents belong to the same semantic level.
  4. 4.
    Distance between P31 and P279 targets: In cases where an entity is simultaneously linked via both P31 and P279, we compute the semantic distance between the target nodes of these two predicates. A large distance may signal that the entity’s classification straddles both instance and subclass semantics in conflicting ways.
By aggregating these four indicators, we produce a risk score that reflects the probability of an entity being semantically inconsistent or structurally ambiguous. This score is then surfaced to the user as a diagnostic aid to determine whether further editorial intervention is warranted.
@@
Fig. 2: Four evaluation dimensions for entity-level semantic risk.
3.3 Semantic Drift Detection Methodology
@@
While the graph-structural risk metrics in the second stage provide in-depth diagnostics, their computational complexity makes them difficult to apply at the scale of the entire Wikidata. To address this scalability challenge, we introduce the third-stage semantic drift detection. This method can be regarded as an efficient proxy for the core ideas of the second stage: by leveraging textual semantic similarity, it approximates the risks arising from connections to semantically distant parent classes (corresponding to dimensions 2 and 4 of the second stage), thereby enabling rapid, large-scale scanning of the entire knowledge graph.
@@
3.3.1 Structural Pre-screening
@@
3.3.2 Text Embedding and Semantic Representation
@@
3.3.3 Semantic Drift Computation

### v2

@@ Abstract.
Abstract.
Wikidata, Data Quality
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
Established in 2012, Wikidata has evolved into one of the internet’s most prominent sources of structured information, encompassing over 120 million entities (Wikidata). Its success stems from a unique model that combines community contributions with data integration from diverse external sources like Gene Ontology and Wikipedia (SHENOY2022100679). This openness has positioned Wikidata as a central resource for knowledge graph research, yet its relatively loose editorial policy introduces significant challenges to its taxonomic integrity and data quality.
While numerous quality issues exist, such as property constraint violations and entity duplication (SHENOY2022100679), a particularly fundamental problem lies within its classification hierarchy. The predicates P31 (instance of) and P279 (subclass of) form the taxonomic backbone of the knowledge graph, connecting over 99% of its entities. However, the misuse and conflation of these two critical relationships have led to what Dadalto et al. (2022) described as a ”large-scale conceptual disarray” (Dadalto2024EvidenceOL). Their work revealed frequent anti-patterns where entities are simultaneously treated as both instances and classes, undermining the logical consistency of the classification system.
The consequences of such taxonomic inconsistencies are far-reaching. A clear distinction between instances and classes is crucial for the performance of downstream applications, including knowledge graph embedding, where it enables more precise modeling of relational patterns (lv-etal-2018-differentiating). Furthermore, structural issues like redundant connections and classification errors can introduce noise that degrades the efficiency and accuracy of reasoning tasks (Wang2023RiverON). For a large, multi-domain knowledge graph like Wikidata, maintaining internal structural coherence is therefore essential for its reliability and utility.
@@ 1. Introduction
2. Related Work
2.1. Wikidata Quality
@@
Wikidata’s status as a premier large-scale open knowledge graph is built upon its principles of community control, plurality, and reliance on secondary data(Wikidata), which have fueled its expansion to over 120 million entities by 2025. However, this scale and openness also introduce significant data quality challenges. A comprehensive study by Shenoy et al. (2021)(SHENOY2022100679) identified a spectrum of quality issues, including property constraint violations, entity duplication, and inconsistent naming conventions. Among these, the conceptual confusion between classes (types) and instances (individuals) stands out as a fundamental and pervasive problem.
@@
Dadalto et al. (2022) (Dadalto2024EvidenceOL)provide compelling evidence of this ”large-scale conceptual disarray” within Wikidata’s multi-level taxonomies. Through an analysis of anti-patterns involving the P31 (instance of) and P279 (subclass of) relations, their work reveals that a vast number of entities are simultaneously treated as both types and instances. This issue is particularly acute in domains such as biology and biochemistry, where concepts like gene and protein are frequently misrepresented. The authors attribute this problem to several root causes: the inherent ambiguity of natural language constructs like ”is-a” (Brachman, 1983(IS-A)), the multi-faceted nature of many real-world concepts, and practical issues such as large-scale automated data imports and the limited adoption of formal classification schemes. These findings underscore the critical need for robust methods to detect and resolve such taxonomic inconsistencies, which is a primary motivation for our current research.”
Patel-Schneider & Doğan (2024) (patelschneider2024classorderdisorderwikidata)conducted a large-scale SPARQL-based analysis of class-order violations in Wikidata, revealing pervasive subclass/superclass misplacements and other hierarchy disorders. They demonstrated that many of these issues can be corrected via manual curation, and provided recommendations for tooling and community practices to improve ontology consistency. This work highlights that taxonomic problems in Wikidata are systemic and repairable.
2.2. Redundant Connections and KG-based tasks
@@
In the context of large-scale knowledge graphs such as Wikidata, the issue of edge redundancy presents a nuanced challenge with both detrimental and beneficial aspects. From a data quality and maintenance perspective, redundancy is traditionally viewed as a negative attribute. Foundational database theories, as discussed by Ferrarotti et al. (2009)(Ferrarotti2009FirstOrderTA), highlight that redundant relations can lead to update anomalies and data inconsistencies. This perspective is echoed in knowledge graph construction, where methodologies often aim to build low-redundant and high-accuracy graphs, treating redundancy as noise that compromises data integrity (Li et al., 2023(low-redun)). Furthermore, formal validation approaches, such as formalizing Wikidata’s property constraints using SHACL and SPARQL, implicitly target redundancy, as redundant edges that are not properly maintained can lead to constraint violations and complicate data validation efforts (Ferranti et al., 2024(Ferranti2024FormalizingAV)).
@@
The negative impact of redundancy is particularly pronounced in downstream knowledge graph reasoning tasks. Research by Wang et al. (2023)(Wang2023RiverON) demonstrates that path redundancy in Graph Neural Network (GNN)-based models can increase transformation error entropy, thereby degrading both the efficiency and accuracy of the reasoning process. Their proposed solution involves a mechanism to prune redundant paths during reasoning, underscoring the performance cost associated with such redundancy.
@@
However, a complete removal of all logically redundant edges may not be the optimal strategy. The utility of a knowledge edge is often context-dependent. Yan et al. (2020)(Yan2020LearningCK) propose a model that learns to generate and filter contextualized knowledge structures for specific commonsense reasoning tasks. Their work suggests that an explicitly stated edge, even if logically inferable and thus ”redundant”, can serve as a powerful and direct signal or a ”shortcut” for reasoning. Eliminating such edges globally could remove valuable heuristics and harm the performance on certain tasks. This indicates that a more sophisticated approach than outright deletion is required. The existing literature thus points towards the need for a filtering mechanism, rather than complete removal, to manage redundancy—a perspective that motivates our present work.
3. Overall Framework
@@
In our research, we conducted a three-stage analysis of the overall data structure of Wikidata to identify and mitigate semantic ambiguities across its classification system.
@@
To support this, we constructed a graph-structured representation of Wikidata using the complete set of P31 and P279 triples. Based on the connectivity patterns among parent classes, intra-graph distances, and hierarchy topology, we defined three dimensions of risk to assess the consistency of each entity’s classification. The resulting scores were integrated into a web-based user interface that enables real-time inspection of the semantic risk associated with any given entity.
@@
4. CME Detection
4.1. Methodology
In Wikidata, the mixed use of P31 and P279 can lead to confusion in classification hierarchy and meaning. The Figure 1 shows some typical errors, as well as the link forms we consider correct in this study. An entity that is linked to a higher level through the P279 relationship should be considered a ”class” entity. As a class entity, it can be an instance of the ”class” (abstract definition) in Wikidata, but it should not be an instance of other entities with actual meaning. Conversely, if an entity is already linked to a meaningful class entity through P31, it should itself be an instance rather than a subclass, and it should not have P279 relationships from other entities. The ’complete and correct’ patterns depicted in Figure 1 are derived from foundational principles of multi-level modeling theory(Dadalto2024EvidenceOL), which enforce a strict stratification between instances (0-order entities) and classes (first-order, second-order, etc.). Any deviation from these patterns, such as an entity being both an instance and a subclass, violates these principles.
Figure 1. Error Types in P31/P279 Relations
@@ 4.1. Methodology
4.2. Results and Evaluation
We selected six representative knowledge domains and performed five rounds of random sampling. In each round, 100 entities were drawn from each domain, resulting in multiple sample sets for evaluation. For every sampled entity, we examined the graph structure formed through its associated P31 (instance of) and P279 (subclass of) relations, and recorded the number of entities that were ultimately identified as structurally incorrect.
The results confirm observations from prior work: in most knowledge domains, structural inconsistencies caused by the misuse of P31 and P279 are widespread, affecting at least 40% of sampled entities. In certain domains—particularly those involving highly technical or hierarchically complex knowledge such as genetics or geography—the proportion of incorrect classifications is significantly higher, reaching 80% to nearly 100%.The results of our random tests are consistent with previous research findings, which proves that our method in this phase is effective.
Figure 2. Error Count Across Five Sampling Rounds in Six Knowledge Domains
In order to first gain an understanding of Wikidata as a whole, we analyzed the all entities with respect to P279 (subclass of) and P31(instance of): There are 4.92 million entities in Wikidata that are used as classes (linked to other entities through the P279 relationship).We define a clean class entity as a class that has exactly one P279 (subclass of) relation and does not participate in any P31 (instance of) relation as a subject, i.e., it does not use P31 to point to any real-world entity class. After excluding class entities that are not referenced by any instances, we obtain 43,065 entities that satisfy our definition of clean classes. For each of these clean classes, we trace their P279 paths upward to identify their corresponding root classes. Classes that share the same root are grouped together, forming independent knowledge-domain-specific entity trees.
Through this process, the 43,065 clean classes are organized into 24,793 distinct knowledge domain trees. Together, these trees contain approximately 10 million instances, accounting for about 10% of all Wikidata entities.
Table 1. Overview of Class Hierarchy Characteristics in Wikidata
Category Count Proportion in Wikidata
Class Entities 4,925,023 4.3%
Clean Classes Entities(have instances) 43,065 0.03%
Hierarchy Root Clean Classes Entities 23,767 0.16%
Instances in Clean Classes Entities 10,458,867 9.2%
All Entities in Wikidata Dump 113,578,686 100.0%
Figure 3. Clean Classes with Instances Number
The figure3 illustrates the distribution of instances across root class entities in Wikidata. Each labeled block represents a root class entity(escape ”Others”), the corresponding number indicates the total number of instances subsumed under that class.
The explicitly shown root classes span a diverse range of knowledge domains, including biological taxonomy (e.g., taxon), astronomy and astrophysics (e.g., astronomical radio source, variable star, high proper-motion star), cultural and creative works (e.g., album, literary work, musical group), built environments and institutions (e.g., building, primary school, branch post office), as well as legal and administrative categories (e.g., United States patent, UK Statutory Instrument).
Figure 4. Root Entity Rank by Instances Numbers
To quantify the structural distribution hidden within the “Others” category, Figure 4 plots the instance count against the root entity rank on a semi-logarithmic scale. It reveals a distinctive heavy-tailed distribution characteristic in Wikidata. While the head of the curve exhibits a precipitous drop—confirming that the vast majority of data volume is concentrated in a handful of dominant super-classes—the tail extends significantly with a gradual decay. the step-like artifacts observed at the far right of the spectrum (ranks >10,000>10,000) represent a vast number of niche root classes with minimal instance counts (single digits), highlighting the extreme sparsity and semantic diversity of the taxonomy’s fringe.
5. Entity-Level Semantic Risk Evaluation
@@
Figure 5. Three Evaluation Dimensions for Entity-Level Semantic Risk.
To assess the degree of semantic ambiguity or multi-meaning risk associated with individual Wikidata entities, we designed a scoring system based on the following three Dimensions:
  1. (1)
    Connectivity Density (Risk Dimension I):
    This dimension quantifies the risk of ambiguity derived from the raw, unpruned dataset. Statistics show that the average number of P31 and P279 relations per entity are approximately 1.068 and 1.001, respectively. An excessively high count suggests that the entity is subject playing multiple conflicting roles or being mapped to overlapping categories.
    We calculate the raw connectivity count $N_{raw}=|P31|+|P279|$ and apply Min-Max normalization to map it to the interval $[0,1]$.
    $\text{Risk}_{\text{I}}(e)=\frac{N_{raw}(e)-\min(N)}{\max(N)-\min(N)}$
    In our data The Max number of P31 and P279 is 59, and the Min number is 1. A higher score directly indicates greater semantic complexity and higher classification risk.
  2. (2)
    Structural Coherence (Risk Dimension II):
    This metric measures the extent to which the entity’s parent concepts are disjointed or topologically scattered. It is calculated based on the cleaned parent set $U^{\prime}(e)$ (after removing direct hierarchical redundancies).
    We first define the connectivity $\mathrm{conn}_{H}(c_{i},c_{j})$ as 1 if the distance $d_{G}(c_{i},c_{j})\leq 3$, and 0 otherwise. To capture risk, we define the Structural Coherence score as the complement of the coherence ratio:
    $\text{Risk}_{\text{II}}(e)=1-\frac{\sum\limits_{c_{i},c_{j}\in U^{\prime}(e),i<j}\mathrm{conn}_{H}(c_{i},c_{j})}{\binom{|U^{\prime}(e)|}{2}}$.
    (For $|U^{\prime}(e)|<2$, we define $\text{Risk}_{\text{II}}(e)=0$ as there is no Structural Coherence risk).
    Under this formulation, a high score (close to 1) implies that the parent classes are isolated from each other (semantic disjointness), signaling a high risk of concept drift or inconsistent classification. Conversely, a score of 0 indicates a perfectly cohesive cluster.
  3. (3)
    Depth Consistency (Risk Dimension III):
    This Dimension evaluates the risk associated with inconsistent abstraction levels within the cleaned set $U^{\prime}(e)$. We calculate the variance of the shortest path lengths to the root node (Q35120):
    $\text{Risk}_{\text{III}}(e)=\mathrm{Var}_{\text{depth}}(e)=\frac{1}{|U^{\prime}(e)|}\sum_{c_{i}\in U^{\prime}(e)}\left(l(c_{i})-\bar{l}\right)^{2}.$
    Since variance measures the spread of data, a higher value inherently indicates that the entity is simultaneously linked to concepts at vastly different granularities (e.g., mixing root-level categories with leaf-level instances). This high disparity reflects taxonomic instability, representing a higher risk of structural error.
The final Composite Risk Score ($S_{risk}$) is computed as a direct weighted sum:
$S_{risk}(e)=w_{1}\cdot\text{Risk}_{\text{I}}(e)+w_{2}\cdot\text{Risk}_{\text{II}}(e)+w_{3}\cdot\text{Risk}_{\text{III}}(e)$
A higher $S_{risk}$ alerts the system to potentially erroneous or low-quality entity entries.
By aggregating these three metrics, we produce a risk score that reflects the probability of an entity being semantically inconsistent or structurally ambiguous. This score is then surfaced to the user as a diagnostic aid to determine whether further editorial intervention is warranted.
6. Semantic Drift Detection Methodology
@@
While the graph-structural risk metrics in the second stage provide in-depth diagnostics, their computational complexity makes them difficult to apply at the scale of the entire Wikidata. To address this scalability challenge, we introduce the third-stage semantic drift detection. This method can be regarded as an efficient proxy for the core ideas of the second stage: by leveraging textual semantic similarity, it approximates the risks arising from connections to semantically distant parent classes , thereby enabling rapid, large-scale scanning of the entire knowledge graph.
@@
This study uses text similarity based on pre-trained language models to approximate semantic consistency, a methodology that is directly inherited from the text enhancement graph embedding paradigm pioneered by KG-BERT (Yao et al., 2019)(yao2019kgbertbertknowledgegraph). Yao et al. confirmed that treating triples as text sequences input to a Transformer can capture semantic plausibility that cannot be recognized by structural information alone. Unlike KG-BERT, which focuses on link prediction, we adapt this hypothesis to the anomaly detection task: that is, a valid categorical link should exhibit extremely high text implicit in vector space.
It is worth noting that the term ’semantic drift’ used in this paper needs to be distinguished from the term of the same name in the field of ontological evolution. In classical work such as SemaDrift (Stavropoulos et al., 2019)(STAVROPOULOS201987) and OntoDrift (Capobianco et al., 2021)(capobianco2020ontodrift), semantic drift refers specifically to the evolution of conceptual connotations over time dimensions (i.e., conceptual drift). In contrast, the semantic drift defined in this paper refers to the topological divergence of entity nodes and their taxonomic ancestors in the semantic vector space under a single static snapshot. This divergence measures the ’conceptual stretch’ in the classification hierarchy, i.e. the extent to which a particular instance is incorrectly classified into a parent class whose semantic category is too broad or irrelevant.
6.1. Structural Pre-screening
@@
6.2. Text Embedding and Semantic Representation
@@
6.3. Semantic Drift Computation